OpenStudy (anonymous):

Perform the division and write the answer without using negative exponents. _72R^2S^6T^3 OVER 27r^6s^3t^8

OpenStudy (anonymous):

-72r^2s^6t^3 over 27r^6s^3t^8

OpenStudy (anonymous):

Do it in pieces, I'll start u off: -72/27 = -8/3 and r^2/r^6 = r^-4 = 1/r^4 U can do the s and t.